Grade 12Magical Mathematics[Interesting Approach]

The half life of first order reaction is 20 minutes. What is the time requires for 99℅ completion of the reaction?

Subham das
6 Years agoGrade 12

1 Answer

Saurabh Koranglekar
6 Years ago
Dear student

20 = 0.693/k

k = 0.693/20= 0.03465

0.03465= (2.303/t) log (1/0.01)

t = 133 min approximately

Regards